OpenStudy (anonymous):

Please help, 1 Question :) What is the value of Log4 2? A. 2/3 B. 1/4 C. 1/3 D. 1/2

OpenStudy (irishboy123):

\(\large z= \log_4 2\) \(\large 4^z = 2\) \(\large z = ?????\)

OpenStudy (anonymous):

D?

OpenStudy (loser66):

yup

OpenStudy (anonymous):

Thanks can you help with another question?

OpenStudy (anonymous):

Solve Log(2x+1)=3 A.999 B.500 C. 999/2 D. 1/2

jimthompson5910 (jim_thompson5910):

You can also use the change of base rule \[\Large \log_{4}(2) = \frac{\log(2)}{\log(4)} = \frac{0.30102999566399}{0.60205999132797} = 0.5 = \frac{1}{2}\]

OpenStudy (anonymous):

Thanks^

OpenStudy (bibby):

follow the same property, log implies the base is 10 \(\log _{10} 2x+1=3 \implies 10^3=2x+1\)

OpenStudy (anonymous):

Is it B?

OpenStudy (bibby):

10^3=1000 1000=2x+1 999=2x x=999/2

OpenStudy (anonymous):

oh ok so its A

OpenStudy (bibby):

lol

OpenStudy (anonymous):

sorry haha thank you :)

OpenStudy (anonymous):

C*

OpenStudy (bibby):

yeah, :p